P(2,3) lies on 4x+3y=17. Find coordinates of points farthest from the line which are 5 units from p

Grade:

1 Answers

Arun
6 years ago
let the points which are at 5 distance from the given line are (a,b)

now
(4a + 3b – 17)/ 5 = ± 5
4a + 3b = 42
&
4a + 3b = – 8
now
distance between (2, 3) & (a, b)
and maximize it.
you will get the value a and b.
hope it helps.
In case of any query please feel free to ask.